# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v1

@@ Abstract
It is well established that pseudo-Hermitian quantum mechanics is a trivial extension of regular quantum mechanics. The modified inner-product space defined through the so-called metric operator, turns out to be the most natural way to represent certain phenomena such as those involving balanced gain and loss. However, for composite systems undergoing pseudo-Hermitian evolution, defining the subsystems is generally considered feasible only when the metric operator is chosen to have a tensor product. In this work, we use arguments from algebraic quantum mechanics to show that the subsystems can be well-defined in every metric space – irrespective of whether or not the metric is of tensor product form. This is done by identifying subsystems with a decomposition of the underlying algebra into sub-algebras. In fact, we show that different decompositions of the underlying $C^{*}-$algebra correspond to choosing different equivalence classes of metric operators. We show how each of the subsystems, defined this way, can be tomographically constructed and that these subsystems satisfy the no-signaling principle. Therefore, we put all the choices of the metric on an equal footing.
@@ I Introduction
It is an axiomatic assumption in quantum theory that the Hilbert space of a composite system has a tensor product decomposition. That is to say, a system with two degrees of freedom, represented by the set of variables $m_{1}$ and $m_{2}$ has the wavefunction $\psi(m_{1},m_{2})\in\mathcal{H}$ in the form $\psi(m_{1},m_{2})=\sum_{i}\psi_{i}(m_{1})\otimes\varphi_{i}(m_{2})\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Such a decomposition allows for a description of the sub-system localized to a sub-space of the total Hilbert space, with operations on this sub-system defined as operations on this sub-space. While this axiom has a footing in some fundamental principles of functional analysis and can be argued from deeper laws of nature  [2, 1], in certain situations, it does not hold good. The most celebrated demonstration of such a deviation is the system of indistinguishable particles [3, 4] and systems with infinite degrees of freedom described by quantum field theory [6, 5]. Another example, and the central topic of this work, is the system that evolves under pseudo-Hermitian Hamiltonians [7, 8]. Pseudo-Hermitian operators are the non-Hermitian operators with real eigenvalues. They arise naturally in certain optical settings with equal gain and loss [9, 10]. It was realized that such evolutions can be seen as unitary evolutions in Hilbert spaces with modified inner product structures induced by the metric operator  [12, 11]. In fact, such a description is necessary to avoid fallacious results like the violation of the no-signaling principle  [13]. For composite systems under such evolutions, the partial trace operation is well-defined if the metric operator allows a tensor product decomposition of the Hilbert space. Such choices of the metric operators are often chosen over other metric operators to define the subsystem  [14, 15]. However, it can so happen that for a given Hamiltonian, none of the choices of the metric operator allow for a tensor product structure [16]. In such situations, one must use an operational definition of the subsystem.
@@
In this work, we outline the problems that arise in defining a subsystem under pseudo-Hermitian evolution and resolve them using tools from the algebraic formulation of quantum mechanics. In the literature, the issue of defining subsystems independently of the Hilbert space structure has been explored by decomposing the underlying $C^{*}$-algebra [17]. We use this algebraic basis to define subsystems in non-trivial inner product spaces. This allows us to put all the choices of the metric operator on equal footing. The central argument of this work is that the different decomposition of the underlying algebra amounts to choosing a particular family of inner product Hilbert spaces to represent a system. Consequently, while a pseudo-Hermitian evolution can be represented as unitary evolution in a family of inner product spaces, different choices of the inner product spaces can result in the probing of different sub-system decompositions. Finally, we show that the no-signaling principle is not broken under this description of the sub-system. While thus far the literature has relied on a genuine tensor product structure of the inner product spaces, we show that under any choice of the metric operator, the no-signaling principle holds for a well-defined sub-system.
@@ I.1 Underlying vector space and the metric
The state of a classical system at a given time is given by a set of real numbers $\{m_{1},m_{2},...\}$ which correspond to a point in the configuration space. In other words, a measurement of the system can yield these values simultaneously. Let $m_{i}\in M_{i}$ where $M_{i}$ is a set of possible values the state can take. We call a system bipartite if the underlying configuration space is $M=M_{1}\times M_{2}$. This means that the experimentalist has the means to measure the parameters of $M_{1}$ and $M_{2}$ at the same time. For example, by placing the spin detectors at different points in space, one can measure the spin and the position of a system. A quantum state is a complex-valued continuous function over the configuration space, given by $\psi(m_{1},m_{2},...)$. The space of all continuous functions on a locally compact Hausdaurff space $M$, denoted by $C^{\infty}(M)$, is a normed vector space that is complete with respect to the metric induced by the norm. This defines a Hilbert space. We are generally interested in a subspace $\mathcal{F}(M)\subset C^{\infty}(M)$, which is also complete with respect to the norm. The elements of this space are the function $f(m_{1},m_{2})$ such that $f(m_{1},\cdot)=f_{2}(\cdot)\in\mathcal{F}_{2}(M_{2})$ and $f(\cdot,m_{2})=f_{1}(\cdot)\in\mathcal{F}_{1}(M_{1})$, where $\mathcal{F}_{1}(M_{1})$ and $\mathcal{F}_{2}(M_{2})$ are set of continuous functions over $M_{1}$ and $M_{2}$. Let the set $\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ denote the closure of the set of linear combinations of functions of the form $\varphi(m_{1},m_{2})=f_{1}(m_{1})\times f_{2}(m_{2})$. Then, we have the equality $\mathcal{F}_{1,2}(M_{1}\times M_{2})=\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ if at least one of the sets $\mathcal{F}_{i}(M_{i})$ is a complete normed vector space   [18].
@@
I.2 Summary of non-Euclidean metric space
@@
Given two elements $\psi$ and $\phi$ on a vector space $\mathcal{V}$, define a Euclidean (or trivial) inner product map $\langle\psi|\phi\rangle\ =\bar{\psi}\phi$, where $\bar{\psi}$ is the conjugate transpose of the element $\psi$. With respect to this inner product, one can define a class of non-Euclidean (or non-trivial) inner products facilitated by a Hermitian, invertible operator $G$ called a metric operator: $\langle\psi|\phi\rangle_{G}\ =\bar{\psi}G\phi=\langle\psi|G\phi\rangle$. We therefore have a class of inner product spaces $\mathcal{H}_{G}:=(\mathcal{V},G)$, where we will denote $\mathcal{H}_{\mathbb{1}}$ simply as $\mathcal{H}$. It can be shown that if $\mathcal{H}_{\mathbb{1}}$ is a Hilbert space, then $\mathcal{H}_{G}$ is also a Hilbert space, i.e. complete with respect to the norm induced by their respective inner product map. The issues in defining subsystems in these Hilbert spaces arise since for an underlying vector space $\mathcal{V}_{1}\otimes\mathcal{V}_{2}$, one has the freedom to endow it with a metric operator that is not in the form of $G_{1}\otimes G_{2}$. In such a case, the partial trace operation on $\mathcal{H}_{G}\rightarrow\mathcal{H}_{1}$ is not well-defined.
@@
An observable is defined as an operator with real expectation values. In $\mathcal{H}_{G}$, such operators must satisfy the relation $G\mathcal{O}=\mathcal{O}^{\dagger}G$, called pseudo-Hermiticity  [8]. An operator can be an observable in two different Hilbert spaces since the metric satisfying the pseudo-Hermiticity condition is not unique: if $\exists$ an invertible operator $T$ s.t. $[T,\mathcal{O}]=0$, then $G^{\prime}=T^{\dagger}GT$ is a compatible metric operator, i.e. $G^{\prime}\mathcal{O}=\mathcal{O}^{\dagger}G^{\prime}$.
@@
A pseudo-Hermitian operator in $\mathcal{H}_{G}$ can be mapped to a Hermitian operator in $\mathcal{H}$ via $\mathcal{O}\rightarrow\mathcal{O}_{\eta}=\eta\mathcal{O}\eta^{-1}$, where $\eta$ is the unique positive square root of the metric operator $G$. Note that $\eta:\psi\rightarrow\eta\psi$ can be seen as a unitary transformation between the spaces $\mathcal{H}_{G}$ and $\mathcal{H}$. Furthermore, since $G^{\prime}$ is another possible metric operator, $\mathcal{O}_{\eta^{\prime}}=\eta^{\prime}\mathcal{O}\eta^{\prime-1}$ is another Hermitisization of the observable $\mathcal{O}$. These two Hermitian forms of $\mathcal{O}$ are unitarily equivalent: $\mathcal{O}_{\eta}=U\mathcal{O}_{\eta^{\prime}}U^{\dagger}$, where $U$ is a unitary operator in $\mathcal{H}$ given by $U=\eta^{\prime}T^{-1}\eta^{-1}$[16] (see Fig.  1). It can be shown that the space of all allowed metric operators compatible with Hamiltonian $H$, can be parametrized by the set of real vectors $\vec{\lambda}$[11]. Given $\{\psi_{i}\}$ and $\{\phi_{j}\}$ the left and right eigenvectors of $H$ (given it has a non-degenerate spectrum), the most general metric is given by $G_{\lambda}=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\phi_{i}|$.
@@
Figure 1: Isometric maps between different metric Hilbert spaces. The state $\psi$ in $\mathcal{H}_{G^{\prime}}$ can be mapped to a state in $\mathcal{H}$ using multiple isometric maps. The resulting states are related by a unitary $U$ in $\mathcal{H}$.
@@ I.3 Open-system perspective
Pseudo-Hermitian evolutions in quantum systems have been experimentally exhibited many times in recent years  [19, 20, 21]. This is typically achieved by introducing an interacting potential wherein the amplitude and frequency of the gain and loss of the probability current are carefully controlled to maintain the potential’s PT-symmetry  [22]. That is, the interacting Hamiltonian is invariant under the action of the parity and time reversal operations. Within certain parameters (referred to as the ”unbroken” PT-symmetry regime) the Hamiltonian has a Pseudo-Hermitian form  [23]. Due to the presence of interaction terms in the Hamiltonian, such an evolution is sometimes seen as open system dynamics, even though the effective Hamiltonian has real eigenvalues. A quintessential feature that separates a pseudo-Hermitian evolution from a genuinely non-Hermitian (non-unitary) evolution is the revival of complete information of the system throughout the dynamics, measured as probability values or the information back-flow. It is argued  [24, 25] that such a complete revival is due to the presence of a finite-sized entangled partner in the environment of the system, where the Hermitisization operation $\psi\rightarrow\eta\psi$ is seen as an embedding of the system into a larger Hermitian system. This description, while illuminating, views pseudo-Hermiticity as a non-trivial extension of quantum mechanics. As we will elucidate in the subsequent sections, the pseudo-Hermitian Hamiltonian, although simulated through interaction terms, indeed induces a unitary evolution.
A large class of open system dynamics follows a Lindblad-type master equation, which can be recast as an evolution under an effective non-Hermitian Hamiltonian and a “jump-term”. If the effective Hamiltonian is pseudo-Hermitian, then one can show that the Lindblad equation is simply the von Neumann equation for a redefined pseudo-Hermitian state, and the jump term vanishes (see B for details).
@@ II The C*-algebra
We start with a C*-algebra of operators $\mathcal{A}$ and a positive linear functional called a state $\omega$ (with norm 1) defined on this algebra. The algebra and the state $(\mathcal{A},\omega)$ together represent the statistical data of the given quantum experiment, e.g. $A\in\mathcal{A}$ can represent the measurement operations that can be performed on a system, and number $\omega(A)$ will then represent the result of such operation. A faithful representation of the algebra is an *-isomorphism to the set of bounded operators on a Hilbert space $\pi:\mathcal{A}\rightarrow\pi(\mathcal{A})\subset\mathcal{B}(\mathcal{H})$. This is possible since the set of bounded operators on a complex Hilbert space is also a C*-algebra and the star operation is naturally defined via $\pi(A)^{*}=\pi(A^{*})$. The GNS construction  [26], is the construction of the Hilbert space description of this system, which involves first constructing a $C^{*}$ algebra on which the state is faithful, i.e., has a trivial kernel. This algebra $\mathcal{U}_{\omega}$ is constructed by first identifying the kernel of the state $K_{\omega}=\{X\in\mathcal{A}\,\text{ s.t. }\,\omega(X^{*}X)=0\}$, and “quotienting out” the kernel from the algebra: $\mathcal{U}_{\omega}:=\mathcal{A}/K_{\omega}$. In other words, the elements of $\mathcal{U}_{\omega}$ consist of the equivalence classes $\psi_{A}=\{A+X\,\text{ s.t. }\,X\in K_{\omega}\}$. One can see that the functional $\omega$ is faithful on $\mathcal{U}_{\omega}$. Moreover, $\mathcal{U}_{\omega}$ is a complex vector space, a property one can check by using the fact that $AX\in K_{\omega}\,\forall\,X\in K_{\omega}$, i.e. $K_{\omega}$ is a left ideal of $\mathcal{A}$. In the next step of the GNS construction, the state functional $\omega$ is used to define an inner product on the vector space $\mathcal{U}_{\omega}$ via $\langle\psi_{A}|\psi_{B}\rangle=\omega(A^{*}B)$. With this added structure, $\mathcal{U}_{\omega}$ becomes a pre-Hilbert space which can be completed to form a Hilbert space $\mathcal{H}_{\omega}=\overline{\mathcal{U}}_{\omega}$. The bar over the algebra denotes the completion with respect to the norm induced by the inner product. Therefore, the set of vectors $\{\psi_{A}\}$ is dense in $\mathcal{H}_{\omega}$. If one defines a representation given by the following action on the states: $\pi_{\omega}(A)\psi_{B}=\psi_{AB}$, then one can easily derive the relation $\omega(A^{*}B)=\langle\psi_{\mathbb{1}}|\pi_{\omega}(A^{*}B)\psi_{\mathbb{1}}\rangle$, where $\mathbb{1}$ is the identity element of the algebra $\mathcal{A}$. Therefore, $\psi_{\mathbb{1}}$ is a cyclic vector of the representation $(\mathcal{H}_{\omega},\pi_{\omega})$. This construction gives us a faithful cyclic representation of the algebra $\mathcal{U}_{\omega}$, with the representation space $\mathcal{H}_{\omega}$, the map $\pi_{\omega}$, and the cyclic vector $\psi_{\mathbb{1}}$, such that $\omega(A)=\langle\psi_{\mathbb{1}}|\pi(A)\psi_{\mathbb{1}}\rangle_{\mathcal{H_{\omega}}}$.

### v2

@@ Abstract
This work outlines a consistent method of identifying subsystems in finite-dimensional Hilbert spaces, independent of the underlying inner-product structure. It has been well established that Hilbert spaces with modified inner-product, defined through the so-called metric operator, turn out to be the most natural ways to represent certain phenomena such as those involving balanced gain and loss resulting in pseudo-Hermitian Hamiltonians. For composite systems undergoing pseudo-Hermitian evolution, defining the subsystems is generally considered feasible only when the metric operator is chosen to have a tensor product form so that a partial trace operation can be well defined. In this work, we use arguments from algebraic quantum mechanics to show that the subsystems can be well-defined in every metric space – irrespective of whether or not the metric is of tensor product form. This is done by identifying subsystems with a decomposition of the underlying $C^{*}$-algebra into commuting sub-algebras. We show that different subsystem decompositions correspond to choosing different equivalence classes of the GNS representation. Furthermore, given a form of pseudo-Hermitian Hamiltonian, the choice of the Hamiltonian compatible metric characterizes the subsystem decomposition and as a consequence, the entanglement structure in the system. We clarify how each of the subsystems, defined this way, can be tomographically constructed and that these subsystems satisfy the no-signaling principle. With these results, we put all the choices of the metric operator on an equal footing.
@@ I Introduction
It is an axiomatic assumption in quantum theory that the Hilbert space of a composite system has a tensor product decomposition. That is to say, a system with two degrees of freedom, represented by the set of variables $m_{1}$ and $m_{2}$ has the wavefunction $\psi(m_{1},m_{2})\in\mathcal{H}$ in the form $\psi(m_{1},m_{2})=\sum_{i}\psi_{i}(m_{1})\otimes\varphi_{i}(m_{2})\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Such a decomposition allows for a description of the sub-system localized to a sub-space of the total Hilbert space, with operations on this sub-system defined as operations on this sub-space. While this axiom has a footing in some fundamental principles of functional analysis and can be argued from deeper laws of nature  [2, 1], in certain situations, it does not hold good. The most celebrated demonstration of such a deviation is the system of indistinguishable particles [3, 4] and systems with infinite degrees of freedom described by quantum field theory [6, 5]. Another example, and the central topic of this work, is the system that evolves under pseudo-Hermitian Hamiltonians [7, 8]. Pseudo-Hermitian operators are the non-Hermitian operators with real eigenvalues. They arise naturally in certain optical settings with equal gain and loss [9, 10]. It was realized that such evolutions can be seen as unitary evolutions in Hilbert spaces with modified inner product structures induced by the metric operator  [12, 11]. In fact, such a description is necessary to avoid fallacious results like the violation of the no-signaling principle  [13]. For composite systems under such evolutions, the partial trace operation is well-defined if the metric operator allows a tensor product decomposition of the Hilbert space. Such choices of the metric operators are often chosen over other metric operators to define the subsystem  [14, 15]. However, it can so happen that for a given Hamiltonian, none of the choices of the metric operator allow for a tensor product structure [16]. In such situations, claiming that subsystems cannot be defined would be incorrect, especially if the experimentalist can probe the different degrees of freedom independently of each other. Therefore, to define subsystems in such Hilbert spaces one must use an operational definition of the subsystems.
@@
In this work, we outline the problems that arise in defining a subsystem under pseudo-Hermitian evolution and resolve them using tools from the algebraic formulation of quantum mechanics. In the literature, the issue of defining subsystems independently of the Hilbert space structure has been explored by decomposing the underlying $C^{*}$-algebra [17, 18]. We use this algebraic basis to define subsystems in non-trivial inner product spaces. This allows us to put all the choices of the metric operator on equal footing. The central argument of this work is that the different decomposition of the underlying algebra amounts to choosing a particular family of inner product Hilbert spaces to represent a system. Consequently, while a pseudo-Hermitian evolution can be represented as unitary evolution in a family of inner product spaces, different choices of the inner product spaces can result in the probing of different sub-system decompositions. Finally, we show that the no-signaling principle is not broken under this description of the sub-system. While thus far the literature has relied on a genuine tensor product structure of the inner product spaces, we show that under any choice of the metric operator, the no-signaling principle holds for a well-defined sub-system.
@@ I.1 Underlying vector space and the metric
The “pure” state of a classical system at a given time is given by a set of real numbers $\{m_{1},m_{2},...\}$ which correspond to a point in the configuration space. In other words, a measurement of the system can yield these values simultaneously. Let $m_{i}\in M_{i}$ where $M_{i}$ is a set of possible values the state can take. We call a system bipartite if the underlying configuration space is $M=M_{1}\times M_{2}$. This means that the experimentalist has the means to measure the parameters of $M_{1}$ and $M_{2}$ at the same time. For example, by placing the spin detectors at different points in space, one can measure a system’s spin and position. A quantum state is a complex-valued continuous function over the configuration space, given by $\psi(m_{1},m_{2},...)$. The space of all continuous functions on a locally compact Hausdaurff space $M$, denoted by $C^{\infty}(M)$, is a normed vector space that is complete with respect to the metric induced by the norm. This defines a Hilbert space. We are generally interested in a subspace $\mathcal{F}(M)\subset C^{\infty}(M)$, which is also complete with respect to the norm. The elements of this space are the function $f(m_{1},m_{2})$ such that $f(m_{1},\cdot)=f_{2}(\cdot)\in\mathcal{F}_{2}(M_{2})$ and $f(\cdot,m_{2})=f_{1}(\cdot)\in\mathcal{F}_{1}(M_{1})$, where $\mathcal{F}_{1}(M_{1})$ and $\mathcal{F}_{2}(M_{2})$ are set of continuous functions over $M_{1}$ and $M_{2}$. Let the set $\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ denote the closure of the set of linear combinations of functions of the form $\varphi(m_{1},m_{2})=f_{1}(m_{1})\times f_{2}(m_{2})$. Then, we have the equality $\mathcal{F}_{1,2}(M_{1}\times M_{2})=\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ if at least one of the sets $\mathcal{F}_{i}(M_{i})$ is a complete normed vector space   [19].
@@
I.2 Overview of quantum mechanics in non-Euclidean metric space
@@
Given two elements $\psi$ and $\phi$ on a vector space $\mathcal{V}$, define a Euclidean (or trivial) inner product map $\langle\psi|\phi\rangle\ =\bar{\psi}\phi$, where $\bar{\psi}$ is the conjugate transpose of the element $\psi$. With respect to this inner product, one can define a class of non-Euclidean (or non-trivial) inner products facilitated by a Hermitian, invertible operator $G$ called a metric operator: $\langle\psi|\phi\rangle_{G}\ =\bar{\psi}G\phi=\langle\psi|G\phi\rangle$. We therefore have a class of inner product spaces $\mathcal{H}_{G}:=(\mathcal{V},G)$, where we will denote $\mathcal{H}_{\mathbb{1}}$ simply as $\mathcal{H}$. It can be shown that if $\mathcal{H}_{\mathbb{1}}$ is a Hilbert space, then $\mathcal{H}_{G}$ is also a Hilbert space, i.e. complete with respect to the norm induced by their respective inner product map. The issues in defining subsystems in these Hilbert spaces arise since for an underlying vector space $\mathcal{V}_{1}\otimes\mathcal{V}_{2}$, one has the freedom to endow it with a metric operator that is not in the form of $G_{1}\otimes G_{2}$. In such a case, the partial trace operation on $\mathcal{H}_{G}$ is not well-defined.
@@
An observable is defined as an operator with real expectation values. In $\mathcal{H}_{G}$, such operators must satisfy the relation $G\mathcal{O}=\mathcal{O}^{\dagger}G$, called pseudo-Hermiticity  [8]. An operator can be an observable in two different Hilbert spaces since the metric satisfying the pseudo-Hermiticity condition is not unique. Let $\mathcal{H}_{G^{\prime}}$ be such a Hilbert space. It is characterized by an isometric map $T:\mathcal{H}_{G^{\prime}}\to\mathcal{H}_{G}$ which satisfies the relation $\eta^{\prime}T^{-1}=U\eta$ where $U$ is any unitary map ($U^{\dagger}U=1$). Therefore the isometries $T$ are parameterized by the operator $U$, hence we will denote them by $T_{U}$. The metric operators are related by $T_{U}^{\dagger}GT_{U}=G^{\prime}$, where this relation is independent of $U$. The operator $\mathcal{O}$ is an observable in both metric spaces if and only if $[T_{U},\mathcal{O}]=0$. Given metric operators $G$ and $G^{\prime}$, this relation holds for a unique unitary $U$.
@@
A pseudo-Hermitian operator in $\mathcal{H}_{G}$ can be mapped to a Hermitian operator in $\mathcal{H}$ via $\mathcal{O}\rightarrow\mathcal{O}_{\eta}=\eta\mathcal{O}\eta^{-1}$, where $\eta$ is the unique positive square root of the metric operator $G$. Note that $\eta:\psi\rightarrow\eta\psi$ is an isometry between the spaces $\mathcal{H}_{G}$ and $\mathcal{H}$. Furthermore, since $G^{\prime}$ is another possible metric operator, $\mathcal{O}_{\eta^{\prime}}=\eta^{\prime}\mathcal{O}\eta^{\prime-1}$ is another Hermitisization of the observable $\mathcal{O}$. These two Hermitian forms of $\mathcal{O}$ are unitarily equivalent: $\mathcal{O}_{\eta}=U\mathcal{O}_{\eta^{\prime}}U^{\dagger}$. In fact, one can show that $[T_{U},\mathcal{O}]=0$ if and only if $\mathcal{O}_{\eta}=U\mathcal{O}_{\eta^{\prime}}U^{\dagger}$[16] (see Fig.  1). It can be shown that the space of all allowed metric operators compatible with Hamiltonian $H$, can be parametrized by the set of real vectors $\vec{\lambda}$[11]. Given $\{\psi_{i}\}$ and $\{\phi_{j}\}$ the left and right eigenvectors of $H$ respectively, (given it has a non-degenerate spectrum), the most general metric is given by $G_{\lambda}=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\phi_{i}|$.
@@
Figure 1: Isometric maps between different metric Hilbert spaces. The state $\psi$ in $\mathcal{H}_{G^{\prime}}$ can be mapped to a state in $\mathcal{H}$ using isometric maps $T_{U}$. The resulting states are related by a unitary $U$ in $\mathcal{H}$.
@@ I.3 Open-system perspective
Pseudo-Hermitian evolutions in quantum systems have been experimentally exhibited many times in recent years  [20, 21, 22]. This is typically achieved by introducing an interacting potential wherein the amplitude and frequency of the gain and loss of the probability current are carefully controlled to maintain the potential’s PT-symmetry  [23]. That is, the interacting Hamiltonian is invariant under the action of the parity and time reversal operations. Within certain parameters (referred to as the ”unbroken” PT-symmetry regime) the Hamiltonian has a Pseudo-Hermitian form  [24]. Due to the presence of interaction terms in the Hamiltonian, such an evolution is sometimes seen as open system dynamics, even though the effective Hamiltonian has real eigenvalues. A quintessential feature that separates a pseudo-Hermitian evolution from a genuinely non-Hermitian (non-unitary) evolution is the revival of complete information of the system throughout the dynamics, measured as probability values or the information back-flow. It is argued  [25, 26] that such a complete revival is due to the presence of a finite-sized entangled partner in the environment of the system, where the Hermitisization operation $\psi\rightarrow\eta\psi$ is seen as an embedding of the system into a larger Hermitian system. This description, while illuminating, views pseudo-Hermiticity as a non-trivial extension of quantum mechanics. As we will elucidate in the subsequent sections, the pseudo-Hermitian Hamiltonian, although simulated through interaction terms, induces a unitary evolution.
A large class of open system dynamics follows a Lindblad-type master equation, which can be recast as an evolution under an effective non-Hermitian Hamiltonian and a “jump-term”. Post-selecting on the trajectories without a jump, the evolution of the system can be cast as an evolution under the non-Hermitian Hamiltonian. If the effective Hamiltonian is pseudo-Hermitian, then one can show that the Lindblad equation is simply the von Neumann equation for a redefined pseudo-Hermitian state, and the jump term vanishes (see A for details), effectively reproducing the unitary dynamics.
@@ II The C*-algebra
We start with a C*-algebra of operators $\mathcal{A}$ and a positive linear functional called a state $\omega$ (with norm 1) defined on this algebra. The algebra and the state $(\mathcal{A},\omega)$ together represent the statistical data of the given quantum experiment, e.g. $A\in\mathcal{A}$ can represent the measurements that can be performed on a system, and the number $\omega(A)$ will then represent the result of such operation. A faithful representation of the algebra is an *-isomorphism to the set of bounded operators on a Hilbert space $\pi:\mathcal{A}\rightarrow\pi(\mathcal{A})\subset\mathcal{B}(\mathcal{H})$. This is possible since the set of bounded operators on a complex Hilbert space is also a C*-algebra and the star operation is naturally defined via $\pi(A)^{*}=\pi(A^{*})$. The GNS construction  [27], is the construction of the Hilbert space description of this system, which involves first constructing a $C^{*}$ algebra on which the state is faithful, i.e., has a trivial kernel. This algebra $\mathcal{U}_{\omega}$ is constructed by first identifying the kernel of the state $K_{\omega}=\{X\in\mathcal{A}\,\text{ s.t. }\,\omega(X^{*}X)=0\}$, and “quotienting out” the kernel from the algebra: $\mathcal{U}_{\omega}:=\mathcal{A}/K_{\omega}$. In other words, the elements of $\mathcal{U}_{\omega}$ consist of the equivalence classes $\psi_{A}=\{A+X\,\text{ s.t. }\,X\in K_{\omega}\}$. One can see that the functional $\omega$ is faithful on $\mathcal{U}_{\omega}$. Moreover, $\mathcal{U}_{\omega}$ is a complex vector space, a property one can check by using the fact that $AX\in K_{\omega}\,\forall\,X\in K_{\omega}$, i.e. $K_{\omega}$ is a left ideal of $\mathcal{A}$. In the next step of the GNS construction, the state functional $\omega$ is used to define an inner product on the vector space $\mathcal{U}_{\omega}$ via $\langle\psi_{A}|\psi_{B}\rangle=\omega(A^{*}B)$. With this added structure, $\mathcal{U}_{\omega}$ becomes a pre-Hilbert space which can be completed to form a Hilbert space $\mathcal{H}_{\omega}=\overline{\mathcal{U}}_{\omega}$. The bar over the algebra denotes the completion with respect to the norm induced by the inner product. Therefore, the set of vectors $\{\psi_{A}\}$ is dense in $\mathcal{H}_{\omega}$. If one defines a representation given by the following action on the states: $\pi_{\omega}(A)\psi_{B}=\psi_{AB}$, one can then derive the relation $\omega(A^{*}B)=\langle\psi_{\mathbb{1}}|\pi_{\omega}(A^{*}B)\psi_{\mathbb{1}}\rangle$, where $\mathbb{1}$ is the identity element of the algebra $\mathcal{A}$. Therefore, $\psi_{\mathbb{1}}$ is a cyclic vector of the representation $(\mathcal{H}_{\omega},\pi_{\omega})$. This construction gives us a faithful cyclic representation of the algebra $\mathcal{U}_{\omega}$, with the representation space $\mathcal{H}_{\omega}$, the map $\pi_{\omega}$, and the cyclic vector $\psi_{\mathbb{1}}$, such that $\omega(A)=\langle\psi_{\mathbb{1}}|\pi(A)\psi_{\mathbb{1}}\rangle_{\mathcal{H_{\omega}}}$.